# The Model Openness Framework: Promoting Completeness and Openness for Reproducibility, Transparency, and Usability in Artificial Intelligence

### v5

@@ 5.2 MOF Class Descriptions
Building upon Class III, level II provides model consumers the complete codebase including libraries and tools needed for training, assessing and testing models themselves. Added elements include:
@@ 6 Model Openness Framework Acceptable Licenses
7 Implementing the Framework
@@ 9.2 Out of Scope
The MOF is not designed to solve all issues related to AI and openness, and its effective implementation will rely on the AI community to be transparent and honest in their reporting of the components of the models that they release and the licenses applied to each. The MOF does not intend to address any of the following as they are best addressed through alternative methods and means: AI safety (including bias, fairness, and trustworthiness), performance testing, red-teaming, security and privacy, components related to model serving, and model provenance.
@@ 10 Conclusion
The MOF provides a clear, actionable methodology for assessing and enhancing the openness and completeness of ML models. It comprehensively outlines specific components that should be openly released, including training data, code, model architecture, model parameters, and documentation, among others, as well as with which licenses. This framework gives model producers a roadmap to follow for reproducible and ethical AI development. By adopting open licenses, as prescribed by the MOF, collaboration and community engagement are enabled, allowing the freedom to use, modify, and distribute models and components under the terms of its license. The tiered classification system incentivizes releasing models with increasing levels of completeness. The widespread adoption of the MOF promises to establish completeness and openness as primary criteria alongside the core tenets of responsible AI, ultimately promoting a more transparent and responsible advancement of AI R&D. We encourage model producers to incorporate the framework into their policies to make open science the standard for model distribution, as well as the wider AI community to recognize and reward complete and open distribution of models. Realizing this vision requires a concerted effort by all AI stakeholders, including researchers, developers, institutions, companies, and governments, to embrace both completeness and openness as core tenets. The benefits for science, business, and society make pursuing model transparency worth the challenge. With carefully designed incentives, policies, and community norms, open source and open science ideals can become the norm in AI, rather than the exception. By working together across domains, we can shape AI advancement to be as complete, open, and empowering as possible. The MOF provides practical guidance for this journey towards more trustworthy AI research and development.
Appendix A Model Openness Tool Functionalities
Figure 1: View models in the MOT catalogue
Figure 2: View model classification with MOT
Figure 3: Evaluate models with MOT
Figure 4: Submit model to the MOT catalogue

### v6

@@ 4.2 MOF Class Descriptions
Building upon Class III, Class II provides model consumers the complete codebase including libraries and tools needed for training, assessing and testing models themselves. Added elements include:
@@ 4.3 Hybrid Releases
5 MOF Components
The following defines the 17 components included in the above three-tier classification system of models (see Table 1). They cover the degree of completeness and openness across all aspects of the development process of an ML model, including training data, model architecture, model parameters, evaluation benchmarks, and documentation.
Note that not all components are required for all classes. Each component section below specifies the classes and it applies to and Table 1 lists the components required for each class.
Note that not all components need to be distributed separately, some MAY be combined. For example, evaluation results MAY be included in the research paper, technical report, or model card rather than published as a standalone artifact. This sort of combination SHOULD however be limited to combining component types that are covered by the LICENSE file for that component or the whole distribution.
The key words "MUST", "MUST NOT", "REQUIRED", "SHALL", "SHALL NOT", "SHOULD", "SHOULD NOT", "RECOMMENDED", "MAY", and "OPTIONAL" in this section are to be interpreted as described in RFC2119 [122].
5.1 Model Architecture (III.1)
The model architecture is the core of any ML project. It can include the ML algorithms, neural network layout, connectivity, activations, and other architectural elements. Examples include transformers (e.g., GPT, BERT), convolutional neural networks (CNNs), recurrent neural networks (RNNs), and graph neural networks (GNNs). While the model architecture is often closely tied to the trained model parameters, sharing the architecture alone allows others to understand the structure of the model without necessitating the release of the fully trained model. The model architecture SHOULD be fully described in the research paper, technical report, or model card, and MUST be distributed as open source code under an OSI-approved open source software license that does not limit its usage and derivative works.
5.2 Model Parameters – Final Checkpoints (III.2)
Trained model parameters MUST be released under an open license. In the case of deep learning models, checkpoints from key intermediate stages of training as well as the final optimizer state SHOULD be included. At a minimum the final model parameters and optimizer state (when applicable) MUST be distributed, whether compressed or uncompressed, in a format compatible with popular deep learning frameworks such as TensorFlow, Keras, PyTorch or the framework independent ONNX file format.
To date, model producers have been releasing model parameters (i.e., weights and biases) using an open source software license, such as Apache 2.0 and MIT, even though model parameters are not compatible with such licenses. Since model parameters are in fact data, model parameters SHOULD be distributed under an open data license, like CDLA-Permissive-2.0. Although licenses designed for OSS are permissive and indemnify the developer from liability, open data licenses are better suited to data-specific considerations such as privacy, ethics, and data rights. Most permissive licenses do not refer to data directly and do not address the ability to modify and redistribute model parameters. This gap could result in a legal obligation to any model consumer if the model producer were to implement royalties after widespread adoption of their model. This is a legal gray area that remains untested. The model architecture and model parameters SHOULD be saved independently in different files for distribution, as each one requires a different format-appropriate open license. This separation allows each component to be studied, modified, redistributed, and used independently of the other.
5.3 Technical Report (III.3)
The technical report, which MAY be in the form of a white paper, provides the necessary documentation for the model consumer to understand the performance, usage, and implications of using the model, but not necessarily enough details to reproduce the model and replicate its results. The technical report MUST be included in the distribution or made available on a permanent open access platform such as arXiv, and SHOULD be linked from the model card. The technical report MAY be omitted if a research paper is provided. The technical report MUST be distributed under an open license that SHOULD be appropriate for documentation, ideally CC-BY-4.0 or CC0.
5.4 Evaluation Results (III.4)
Detailed quantitative metrics and qualitative results from evaluating the model MUST be reported. They MAY be included in the technical report, the research paper, or the model card. Tests can evaluate any factor, not limited to model efficiency, accuracy, performance, fairness and bias evaluations, toxicity, truthfulness and so forth. Producers MUST include benchmark test results, whether industry standard benchmarks or custom benchmark tests that were developed. If industry standard benchmark tests or test suites are used, the test suite name, test name and version number MUST be included with the results. If custom benchmarks were developed, whether in code or any form of media including text, images, the custom benchmarks MUST be included in full for validation. The raw outputs of the model evaluation MUST be distributed under an open license that SHOULD be appropriate for content like CC-BY-4.0.
5.5 Model Card (III.5)
A model card provides metrics, usage guidance, and details about a model [21]. Model cards SHOULD cover model details, intended uses, factors, evaluation, risks, and mitigations related to the model. The model card itself MUST be distributed under an open license that SHOULD be appropriate for documentation, ideally CC-BY-4.0.
5.6 Data Card (III.6)
A data card provides summary statistics and key information about a dataset to enhance understanding of its composition [22]. Following guidelines from the Data Nutrition Project [123], data cards SHOULD describe metrics about the features, instances, intended uses, motivation, and collection process. Data cards help identify potential biases in datasets and guide proper usage by downstream usage. They also contribute to reproducibility and transparency by detailing the entire data preparation process. The data card MUST be distributed under an open license that SHOULD be appropriate for documentation, ideally CC-BY-4.0.
5.7 Sample Model Outputs (III.7)
Sample outputs are text samples, images, videos, software code, audio, 3D assets, metadata or any other potential output generated from the model, including predictions and probabilities. Sample outputs are RECOMMENDED but not required. They MAY be omitted. If they are included in the distribution, they MUST be shared publicly without copyright or restrictions where legally permitted so that they can be redistributed with the release. For certain sensitive domains, generated examples MAY be anonymized or simulated if needed. In the event where model outputs are not copyrightable, the outputs SHOULD be released without a license, and this SHOULD be noted in the LICENSE file.
5.8 Training, Validation, and Testing Code (II.2)
The full code for training, validating and testing the model, including model construction, training loop, hyperparameter selection, and checkpointing SHOULD be distributed as open source software. Any fine-tuning code, reinforcement learning code or any other method that otherwise modifies that model parameters or code that implements adapters that ultimately affect the performance of the model MUST be included. Comments explaining the approach SHOULD be included in the code, ideally following PEP 8 style guide for Python code. It is also RECOMMENDED to include with the release the log files generated during training. The training, validation, and testing code itself MUST be released under an OSI-approved open source software license while any logs SHOULD be covered by a permissible open content license like CC-BY-4.0.
5.9 Inference Code (II.3)
The availability of inference code facilitates complete replication of the performance of the model, and it informs the model consumer about how to use the model most effectively for their applications. Code for performing inference includes any data preprocessing or postprocessing required during inference and possibly any model optimizations and dependencies like external libraries. It MUST include any code required to fully replicate the benchmark results for the model. The inference code MUST be released under an OSI-approved open source software license.
5.10 Evaluation Code (II.4)
Evaluation code, evaluation data, and evaluation results are separate components in the MOF. This is due to the fact that some benchmarks are written in code and other benchmarks only use data; for instance, text used to evaluate an LLM or images used to evaluate a computer vision model. Many benchmark tests are a combination of both code and data used to evaluate a model, which includes the scripts needed to load the data and run benchmark tests. Since code and data require different licenses, they are separate components. Depending on the nature of the model and the methods used to evaluate it, the distribution MAY include one or both of evaluation code and data. Any code used for model evaluation and benchmarking MUST be included and distributed under an OSI-approved open source software license.
5.11 Evaluation Data (II.5)
When the model is evaluated with data (be it any media format including text, images, videos, audio, 3D data, and so forth), that evaluation data MUST be included with the distribution. Where the model producer relies on standard benchmark tests that are widely disseminated, they MAY be omitted from the distribution, but they MUST be described in the technical report, research paper, or model card, along with the version of the test. The evaluation data MUST be released under a data or content appropriate open license like CDLA-Permissive-2.0, CC-BY-4.0 or CC0.
5.12 Supporting Libraries and Tools (II.6)
Any supporting code libraries, utilities, or tools developed in the course of the development of the model SHOULD be distributed under an OSI-approved open source software license. This includes data loaders, visualization code, simulation environments, etc. Use of existing and custom open source tools SHOULD also be documented. Any of the following tools and libraries SHOULD also be included:
Presumably most libraries and tools used already have their own licenses, but if the model producer created their own libraries or tools they MUST include them with the distribution under an OSI-approved open source software license.
5.13 Research Paper (I.2)
The research paper detailing the model methodology, results, and analysis MUST be included in the distribution or made available on a permanent open access platform such as arXiv, and SHOULD be linked from the model card. The research paper MUST be released under an open license appropriate for documentation, ideally CC-BY-4.0. The following structure is RECOMMENDED but not required: abstract, introduction, related work, methods, results, discussion, conclusion, references. The research paper MAY be replaced with a detailed technical report providing the same information and distributed under an open license appropriate for documentation, ideally CC-BY-4.0.
5.14 Datasets (I.3)
Data is the lifeblood of ML models and is the most often held back element in the release of a model. Datasets include training data which is data used for any form of model training including pre-training, fine-tuning, alignment using reinforcement learning techniques or data used for other methods that otherwise modify the weights of the model. Datasets also include data used for model validation and testing as well as data used with benchmark tests. The datasets component also includes tokenized datasets when present. Data can be any form or combination of media, whether text, code, images, videos, audio, 3D objects, URIs and any other data used for training, validation and testing purposes. Datasets also include any metadata. This includes anything from annotation data like labels, bounding boxes and key points to attribution, bitrates, resolution and other metadata relevant to a dataset used in the model development process. The datasets used to develop the model MUST be provided, in the public domain, as copyrighted data, or under any form of license. They SHOULD be released under an open license, preferably CC-BY-4.0 or CC-0. Any limits on sharing due to privacy or sensitivity SHOULD be documented. Both pre- and post-processed data SHOULD be supplied, however producers MAY provide instead links to any curated raw datasets online if they are accompanied by data preprocessing code.
5.15 Data Preprocessing Code (I.4)
The data preprocessing code is all code used for preprocessing, cleaning, and formatting the training, validation, and testing data for a model. It also includes code used to transform fine-tuning data and code that is used for alignment tasks like Reinforcement Learning from Human Feedback (RLHF). Other data preprocessing code such as code for data ingestion when appropriate, feature engineering, data augmentation and tokenization is also included. The data preprocessing code MUST be released using an OSI-approved open source software license.
5.16 Model Parameters – Intermediate Checkpoints (I.5)
In addition to the final checkpoints and optimizer states, for Class I models, the checkpoints and optimizer states (when applicable) from key intermediate stages of training along with the log files MUST be included and distributed under an open license. Intermediate model parameters SHOULD be distributed under an open data license, such as CDLA-Permissive-2.0.
5.17 Model Metadata (I.6)
There are other forms of metadata that can provide additional context about the model, such as the version of the framework used to create it and custom tags or descriptions provided by the developer including model and data lineage information. There is no particular requirement or profile for this type of metadata and it can reveal anything the developer would like to include with the shipped model. This information can help with model management, especially when working with multiple versions of models or conducting experiments. Often the metadata is exported from or loaded by a metadata store. The model metadata MAY be included in the model card, research paper, or technical report. Any model metadata SHOULD be covered by an open data license such as CDLA-Permissive-2.0.
5.18 Model Openness Configuration File
The MOF configuration file MUST be included in any distribution. It describes what model components are included in the release and what license covers each component. The file itself MUST be distributed under an open license and SHOULD be distributed under the CC-BY-4.0 license.
@@ 6 Model Openness Framework Acceptable Licenses
7 Adopting the Model Openness Framework
@@ 9.2 Out of Scope
The MOF is not designed to solve all issues related to AI and openness, and its effective adoption will rely on the AI community to be transparent and honest in their reporting of the components of the models that they release and the licenses applied to each. The MOF does not intend to address any of the following as they are best addressed through alternative methods and means: AI safety (including bias, fairness, and trustworthiness), performance testing, red-teaming, security and privacy, components related to model serving, and model provenance.
@@ 10 Conclusion
The MOF provides a clear methodology for evaluating and enhancing the openness and completeness of ML models. It outlines specific components that should be openly released, including training data, code, model architecture, model parameters, and documentation, among others, as well as with which licenses. This framework gives model producers a roadmap to follow for reproducible and transparent AI development. By adopting open licenses, as prescribed by the MOF, we can foster collaboration and community engagement, allowing the freedom to use, study, modify, and distribute models and components under the terms of its license. Furthermore, the tiered classification system incentivizes releasing models with increasing levels of completeness. The widespread adoption of the MOF promises to establish completeness and openness as core tenets of responsible AI, ultimately promoting a more transparent and trustworthy advancement of AI R&D. We encourage model producers to incorporate the framework into their policies to make open science the standard for model distribution, as well as the wider AI community to recognize and reward complete and open distribution of models. Realizing this vision requires a concerted effort by diverse AI stakeholders to embrace both completeness and openness as core tenets of AI R&D. However, the benefits for science, innovation, and society make pursuing model openness worth the challenge. With carefully designed incentives, policies, and community norms, open source and open science ideals can become the norm in AI R&D, rather than the exception.
Appendix A Functionalities of the Model Openness Tool
Figure 1: View models in the catalogue of the Model Openness Tool
Figure 2: View a model’s classification with the Model Openness Tool
Figure 3: Evaluate models with the Model Openness Tool
Figure 4: Submit models to the model catalogue with the Model Openness Tool